OpenStudy (anonymous):

int_{inf}^{?inf} 1/(1+x ^{4}) dx

sam (.sam.):

What? lol can you rewrite it in the comments?

OpenStudy (anonymous):

I"m assuming its an integral, remember: \[\int\limits_{-\infty}^{\infty} f(x) dx = \int\limits_{-\infty}^{0} f(x) dx + \int\limits_{0}^{\infty} f(x) dx\] = \[\lim_{t \rightarrow -\infty}\int\limits\limits_{-t}^{0} f(x) dx + \lim_{t \rightarrow \infty}\int\limits\limits_{0}^{t} f(x) dx\]

OpenStudy (dumbcow):

you have to use partial fractions, but the algebra gets messy http://www.wolframalpha.com/input/?i=integrate+1%2F%281%2Bx^4%29+dx

OpenStudy (anonymous):

\[\int\limits_{\inf}^{\inf}1/(1+x ^{4}) dx\]

OpenStudy (anonymous):

@dumbcow gud answer.

OpenStudy (anonymous):

@.Sam. I commented the question. You can try.